The hypotenuse of a right triangle is the square root of 13cm. One leg is 1 cm shorter then twice the length of the other leg. Find the lengths of the legs of the right triangle.
Answer by mananth(16949)

let the length of one leg be x cm
the other leg = 2x-1 cm
Hypotenuse = sqrt(13)
...
x^2+(2x-1)^2=(13)
x^2+4x^2-4x+1=13
5x^2-4x+1-13
5x^2-4x-12=0
5x^2-10x+6x-12=0
5x(x-2)+6(x-2)=0
(x-2)(5x+6)=0
x= 2
...
length of one leg = 2cm
other leg = 2x-1 = 3 cm
...
m.ananth@hotmail.ca